tim is 3 times as old as tina. 3 years ago, the difference in their age was 10. How old are they now?
Answer by alicealc(293)

Tim's age = m
Tina's age = n

m = 3*n
3 years ago:
(m - 3) - (n - 3) = 10
3n - 3 - n + 3 = 10
2n = 10
n = 10/2 = 5
m = 3*5 = 15

so, Tim is 15 years old now and Tina is 5 years old now